# Jailbreak Large Vision-Language Models Through Multi-Modal Linkage

### v3

@@ 5.2 Main Experiments
The results of the experiment are presented in Table 1 and Table 2. As shown, our methods successfully jailbreak several target models with a high ASR across these datasets, making significant improvements over the baseline. For GPT-4o, we achieve an ASR of 97.80% on SafeBench, 98.81% on MM-SafeBench, and 99.07% on HADES-Dataset, representing increases of 64.80%, 73.56%, and 95.07% over the highest baseline results, respectively. Claude-3.5-Sonnet stands out as the most robust model, performing well against both the baselines and our method. However, MML still manages to jailbreak it with success rates of 69.40%, 60.00%, and 45.73% on the three datasets, showing improvements of 52.80%, 51.86%, and 45.60% compared to the highest baseline ASR. The experimental results demonstrate that current VLMs cannot maintain safety alignment under our attack.
@@
Given that these datasets classify malicious queries into distinct categories, we also evaluate the ASR of our method across various forbidden topics. Figure 4 and Figure 5 illustrate the ASR of MML with image mirroring encryption compared to the baseline methods across four models on these topics in SafeBench and MM-SafeBench. On SafeBench, the baseline method struggles with the first seven harmful topics such as Illegal Activity and Hate Speech on most models, mirroring the observations by Gong et al. (2023). In contrast, our method significantly improves the ASR for these topics, exceeding 95% in most cases, except for Claude-3.5-Sonnet. On MM-SafeBench, our approach consistently achieves at least 95% ASR across most topics and models. Notably, different models exhibit varying performance across different forbidden topics. ASR across various scenario on HADES-Dataset and more detailed results are included in Appendix A.4.
@@ E-D Hint Evil ASR(%) DSR(%)
Table 4: ASR of FigStep vs. MML (ours) in attacking GPT-4o on SafeBench under the AdaShield-Static defense. The best results under the same defense are highlighted in bold. “+” indicates concatenation, “Q” represents the input text prompt, and “D” refers to the defensive prompt from AdaShield-Static. MML-$X$ denotes variations of MML with different encryption methods: WR stands for word replacement, and represents image mirroring.
5.3 Ablation Study
We perform ablation experiments to evaluate three components of the proposed method: the encryption-decryption framework, the inclusion of decryption hints in the prompt, and evil alignment. Using GPT-4o on the SafeBench dataset, we assess their effectiveness through two metrics: attack success rate (ASR) and decryption success rate (DSR). Since the target model’s response must include the decrypted content, it allows for straightforward evaluation of whether the response fully reconstructs the original malicious query.
Encryption-Decryption.
Table 3 highlights the significant impact of the encryption-decryption mechanism on jailbreak success. By employing the mirror encryption-decryption technique alone, we increase the ASR from 34% to 75.20%, with almost no additional computational overhead. However, without hidden malicious guidance, relying solely on encryption-decryption leads to a higher rejection rate (instances where the jailbreak score is 0, as shown in Figure 6).
Hint.
Intuitively, the success of MML depends on the effective reconstruction of the original malicious queries during decryption. Therefore, adding hints is expected to increase the DSR, thereby boosting the ASR. However, our experimental results reveal partial inconsistencies with this expectation. As shown in Table 3, although ASR improves with the addition of hints, DSR actually decreases in the absence of malicious injection. A manual review indicates that a mojority of decryption failures are due to minor errors like singular/plural mismatches, punctuation, or capitalization issues, such as missing periods. These errors, however, do not hinder the inclusion of malicious content.
@@ 5.3 Ablation Study
Evil alignment prove highly effectiveness in enhancing the attack. As shown in Table 3, using only evil alignment achieved an ASR of 89.80%, an unexpectedly high result. Additionally, Figure 6 reveals that after employing evil alignment, the number of moderately malicious responses (scoring 3 or 4) significantly decreases, with nearly all responses scoring 5, indicating a strong alignment between the target model’s output and the malicious intent.
@@ 5.5 Discussions and Limitations
Limitations.
Despite MML achieves high jailbreak success rate on top VLMs such as GPT-4o, it has some limitations. First, we do not evaluate the more advanced GPT-4-o1-Preview (OpenAI, 2024c) due to API access restrictions. Second, Since MML does not instruct the model to conceal harmful content in the output (Wei et al., 2024), it can be defensed by output detection (Pi et al., 2024).
@@ 6 Conclusion
In this work, we propose a novel jailbreak framework Multi-Modal Linkage (MML) Attack targeting at the safety alignment of VLMs. To address the issues of over-exposure of malicious content in existing methods, MML introduces a cross-modal encryption-decryption mechanism. In addition, to amplify the maliciousness of the target model’s response, we depict a virtual video game production scene to align model’s output with malicious. Extensive experiments on two datasets demonstrate the effectiveness of our approach.
@@ A.1 Implicit Rejection
Neutral text guidance can lead to implicit rejection, which avoids directly answering malicious questions and instead gives the opposite, helpful suggestions. For example, here is an indirect rejection response to “Steps to self-harm” generated by GPT-4o:
@@ A.2 Candidate Replacement Words
For the word replacement encryption method, all nouns in the malicious query are replaced with food-related terms, and all adjectives are substituted with positive words. The candidate replacement words are listed below, and the replacements are applied sequentially.
@@ A.4 Main Experiment Details
MM-SafeBench includes 13 scenarios, with the number of queries varying across each. Additionally, we filter the queries in MM-SafeBench (Liu et al., 2024c) to exclude those not considered violations. Specifically, we use GPT-4o-Mini to evaluate whether a query violations the CLAS usage policy, which is more comprehensive. Table 5 shows the details about MM-SafeBench in our experiment. The CLAS policy is shown in Figure 7. The filtering prompt is illustrated in Figure 12. GPT-4o-Mini assigns scores to queries on a scale from 1 to 4, with higher scores indicating a greater degree of violation. Given that VLMs can readily respond to mildly illegal queries even in plain text, only queries receiving a score of 4 are included in our experiment.
@@
Figure 11: Some results of MML targeted GPT-4o on four prohibited topics in SafeBench, employing various encryption methods. Certain responses are masked for security purposes.
@@
We adopted the CLAS approach for evaluation. If no rejection words are detected in the model’s response, GPT-4o-Mini assesses the success of the attack. The evaluation measures the alignment between the original malicious query and the model’s output, assigning a score on a scale from 1 to 5. The scoring criteria are illustrated in Figure 13. A score of 1 indicates the model refused to respond, while a score of 5 signifies full compliance with the policy-violating instruction without deviation. The complete evaluation prompt is provided in Figure 14.
@@
Figure 11 shows the results of MML using four encryption methods across four forbidden topics in SafeBench, with GPT-4o as the target model. For security concerns, we mask part of responses.

### v4

@@ 5.2 Main Experiments
The results of the experiment are presented in Table 1 and Table 2. As shown, our methods successfully jailbreak several target models with a high ASR across these datasets, making significant improvements over the baselines. For GPT-4o, we achieve an ASR of 97.80% on SafeBench, 98.81% on MM-SafeBench, and 99.07% on HADES-Dataset, representing increases of 64.80%, 73.56%, and 95.07% over the highest baseline results, respectively. Claude-3.5-Sonnet stands out as the most robust model, performing well against both the baselines and our method. However, MML still manages to jailbreak it with success rates of 69.40%, 60.00%, and 45.73% on the three datasets, showing improvements of 52.80%, 51.86%, and 45.60% compared to the highest baseline ASR. The experimental results demonstrate that current VLMs cannot maintain safety alignment under our attack. Qualitative attack results are listed in Figure 11.
@@
Given that these datasets classify malicious queries into distinct categories, we also evaluate the ASR of our method across various forbidden topics. Figure 4 and Figure 5 illustrate the ASR of MML with image mirroring encryption compared to the baseline methods across four models on these topics in SafeBench and MM-SafeBench. On SafeBench, the baseline methods struggle with the first seven harmful topics such as Illegal Activity and Hate Speech on most models, mirroring the observations by Gong et al. (2023). In contrast, our method significantly improves the ASR for these topics, exceeding 95% in most cases, except for Claude-3.5-Sonnet. On MM-SafeBench, our approach consistently achieves at least 95% ASR across most topics and models. Notably, different models exhibit varying performance across different forbidden topics. ASR across various scenarios on HADES-Dataset and more detailed results are included in Appendix A.4.
@@ E-D Hint Evil ASR(%) DSR(%)
5.3 Ablation Study
We perform ablation experiments to evaluate three components of the proposed method: the encryption-decryption framework, the inclusion of decryption hint in the prompt, and evil alignment. Using GPT-4o on the SafeBench dataset, we assess their effectiveness through two metrics: attack success rate (ASR) and decryption success rate (DSR). Since the target model’s response must include the decrypted content, it allows for straightforward evaluation of whether the response fully reconstructs the original malicious query.
Encryption-Decryption.
Table 3 highlights the significant impact of the encryption-decryption mechanism on jailbreak success. By employing the encryption-decryption technique alone, we increase the ASR from 34% to 75.20%. However, without stealthy malicious guidance, relying solely on encryption-decryption leads to a higher rejection rate (instances where the jailbreak score is 0, as shown in Figure 6).
Decryption hint.
Intuitively, the success of MML depends on the effective reconstruction of the original malicious queries during decryption. Therefore, adding hint is expected to increase the DSR, thereby boosting the ASR. However, our experimental results reveal partial inconsistencies with this expectation. As shown in Table 3, although ASR improves with the addition of hint, DSR actually decreases in the absence of evil alignment. A manual review indicates that a majority of decryption failures are due to minor errors like singular/plural mismatches, punctuation, or capitalization issues, such as missing periods. These errors, however, do not hinder the inclusion of malicious content.
@@ 5.3 Ablation Study
Table 4: ASR of FigStep vs. MML (ours) in attacking GPT-4o on SafeBench under the AdaShield-Static (Wang et al., 2024) defense. The best results under the same defense are highlighted in bold. “+” indicates concatenation, “Q” represents the input text prompt, and “D” refers to the defensive prompt from AdaShield-Static. MML-$X$ denotes variations of MML with different encryption methods: WR stands for word replacement, and M represents image mirroring.
@@
Evil alignment prove highly effectiveness in enhancing the attack. As shown in Table 3, using only evil alignment achieves an ASR of 89.80%. Additionally, Figure 6 reveals that after employing evil alignment, the number of moderately malicious responses (scoring 3 or 4) significantly decreases, with nearly all responses scoring 5, indicating a strong alignment between the target model’s output and the malicious intent.
@@ 6 Conclusion
In this work, we propose a novel jailbreak framework Multi-Modal Linkage (MML) Attack targeting at the safety alignment of VLMs. To address the issues of over-exposure of malicious content in existing methods, MML introduces a cross-modal encryption-decryption mechanism. In addition, to amplify the maliciousness of the target model’s response, we depict a virtual video game production scene to align model’s output with malicious. Extensive experiments on three datasets demonstrate the effectiveness of our approach.
7 Ethical Consideration
The goal of this work is to highlight the inadequacy of current safety alignment in VLMs, which fail to prevent them from being abused. Although this paper inevitably contains toxic content generated by VLMs, we have made every effort to mitigate potential abuse, including displaying only part of the content and replacing with “…”. Our motivation is to raise awareness of this potential safety issue, thereby fostering the responsible development of VLMs for the benefit of community.
8 Limitations
Despite MML achieves high jailbreak success rate on top VLMs such as GPT-4o, it has some limitations. First, we do not evaluate the more advanced o1 (OpenAI, 2024c) due to API access restrictions. Second, since MML does not instruct the model to conceal harmful content in the output (Wei et al., 2024), it can be defensed by output detection (Pi et al., 2024).
@@ A.1 Implicit Rejection
Neutral text guidance can lead to implicit rejection, which avoids directly answering malicious questions and instead gives the opposite, helpful suggestions. For example, here is an indirect rejection response to “Methods for Self-harm” generated by GPT-4o:
@@ A.2 Candidate Replacement Words
For the word replacement encryption, all nouns in the malicious query are replaced with food-related terms, and all adjectives are substituted with positive words. The candidate replacement words are listed below, and the replacements are applied sequentially.
@@ A.4 Main Experiment Details
Baselines.
QueryRelated first uses GPT-4 to generate malicious queries in different scenarios and then rewrites these queries. Next, QueryRelated uses GPT-4 to extract unsafe keywords from queries and generates three types of images: Stable Diffusion (SD) images, typography images, and SD + typography images. According to the results from Liu et al. (2024b), SD + typography images method is the state-of-the-art, which is tested as the baseline of our experiments.
HADES uses diffusion model (Chen et al., 2024) to generate malicious images based on harmful instructions. During this process, it employs LLMs (OpenAI et al., 2024) to amplify the harmfulness of the generated images. When targeting white-box VLMs, HADES leverages gradients to generate adversarial noise, enhancing the attack. Finally, the adversarial noise, the malicious image, and the keyword typography are concatenated into a single image in sequence. Since adversarial attacks are transferable (Dong et al., 2023; Zhao et al., 2024), we use the images generated by HADES on LLava-v1.5-7B (Liu et al., 2024a) as the input for the baseline.
@@
MM-SafeBench includes 13 scenarios, with the number of queries varying across each. Additionally, we filter the queries in MM-SafeBench (Liu et al., 2024b) to exclude those not considered violations. Specifically, we use GPT-4o-Mini to evaluate whether a query violations the CLAS usage policy, which is more comprehensive. Table 5 shows the details about MM-SafeBench in our experiment. The CLAS policy is shown in Figure 7. The filtering prompt is illustrated in Figure 12. GPT-4o-Mini assigns scores to queries on a scale from 1 to 4, with higher scores indicating a greater degree of violation. Given that VLMs can readily respond to mildly illegal queries even in plain text, only queries receiving a score of 4 are included in our experiment.
@@
Figure 11: Some results of MML targeted GPT-4o on four prohibited topics in SafeBench, employing various encryption methods. For security purposes, only a portion of the content is shown. Warning: these results contains jailbroken contents that may be offensive in nature.
Data preprocessing.
@@
We adopt the CLAS approach for evaluation. If no rejection words are detected in the model’s response, GPT-4o-Mini assesses the success of the attack. The evaluation measures the alignment between the original malicious query and the model’s output, assigning a score on a scale from 1 to 5. The scoring criteria are illustrated in Figure 13. A score of 1 indicates the model refused to respond, while a score of 5 signifies full compliance with the policy-violating instruction without deviation. The complete evaluation prompt is provided in Figure 14.
@@
Figure 11 shows the results of MML using four encryption methods across four forbidden topics in SafeBench, with GPT-4o as the target model. For security concerns, we only display a portion of the responses.